OpenStudy (anonymous):

MEDAL AND FAN In the supermarket, there is a scale that measures the weight of fruit and vegetables. The needle is 24 cm long and it rotated 96° when Charlie placed four tomatoes on it. How far did the needle travel as it rotated?

OpenStudy (anonymous):

It moved 96/360 th of the circumfrence of a circle with diameter of 24*2.

OpenStudy (anonymous):

1152π 2304π 5 pi over 64 64 pi over 5

OpenStudy (anonymous):

WHICH ONE AD WHY?

OpenStudy (anonymous):

\[\frac{96 \pi 48}{360}=\frac{64 \pi }{5}=40.2124 \]

OpenStudy (anonymous):

TAHNK YOU SOO MUCH

OpenStudy (anonymous):

All of these clock type problems are siimilar. Like the hour hand moves 1/12 of the angle that the minute hands moves. Thank you for the medal.